Here is one that i have to solve with the substitution method and determine whether the equations are independent, dependent, or consistent...so lost.
x=y-5 first line
3x-2y=4 second line
why am i not getting it?

Answer by scott8148(6628)

SUSTITUTING the value of x from the first line into the second line
__ 3(y-5)-2y=4 __ 3y-15-2y=4 __ y=19

x=19-5 __ x=14

unique solutions for x and y means independent equations